# MoESD: Unveil Speculative Decoding’s Potential for Accelerating Sparse MoE

### v2

@@ 4 Experiments
We use Qwen2-57B-A14B-Instruct and Qwen2-0.5B-Instruct qwen2 as the target and draft model. The original Qwen2-57B-A14B-Instruct activates 8 experts out of 64. When we need to examine MoEs with different sparsity, we modify the num_experts_per_token in the config.json. We use HumanEval humaneval and MT-bench mtbench datasets for code generation and conversation tasks, as previous works li2024eagle ; hass ; cai2401medusa have done. The sequence lengths of the tokenized prompts of these two datasets range from 38 to 391, and 5 to 356, respectively.
@@
We used the existing vllm kwon2023vllm framework for our experiments to verify theoretical predictions. Vllm supports batched speculative decoding, cudagraph optimization, and reports comprehensive data such as $T_{D}$, $T_{T}$, $T_{reject}$ and $\sigma$, thus being suitable for our experiments. To prevent unstable performance at the beginning, all data were obtained by averaging the results from the last five of ten total runs. We conducted experiments on two types of GPUs, referred to as A and B hereafter. Both GPU A and GPU B has 80GB memory, but their peak computational power and memory bandwidth are different. To accommodate the FP16 57B target model, a minimum of two machines is required for both types of GPUs.
@@ 4.1 Speedup Trend of Speculative Decoding for MoE
Figure 2: (a),(b): SD speedup (left y-axis) as a function of batch size and corresponding target efficiency values (right y-axis). Across different hardware platforms, SD acceleration first increases and then decreases, verifying our theoretical predictions. The target efficiency shows consistent trends with final speedup. (c): Comparison of target efficiency between MoE and dense model.
Table 1: The peak speedup ($\mathbf{x}$) of SD across different datasets, temperatures, $\gamma$s and hardware
@@
2xGPU B humaneval 0.0 15.96 9.34 0.95 1.71 15.96 7.95 0.93 2.01 15.96 6.96 0.90 2.29
@@
Figure 2(a) and 2(b) plot the end-to-end SD speedup (left y-axis) for MoE across various settings, validating our theoretical prediction about acceleration behavior. As batch size grows, speedup initially increases due to expert loading saturation, and then decreases due to compute-boundness. We denote the maximal speedup across batch sizes as $\mathbf{x}$ and summarize the results in Table 1. SD achieves higher acceleration with longer $\gamma$ for tasks with more predictable patterns (e.g., code generation) or less randomness (e.g., lower temperature), aligning with conclusions from previous research. Besides, SD speedup is more significant on GPU B. This is because GPU B has a higher ridge point, providing more arithmetic units for verification. SD is also expected to achieve obvious acceleration on NPUs with high ridge points like Ascend.
Figure 2(a) and  2(b) also highlights the effectiveness of our metric target efficiency. It is computed as $\frac{T_{T}(B,1)}{T_{T}(B,\gamma)}$ as explained in Sec. 3.1, where both $T_{T}(B,1)$ and $T_{T}(B,\gamma)$ are obtained from vllm runtime logs. Target efficiency values are annotated on the right y-axis, showing a consistent trend with the end-to-end speedup. In contrast, the acceptance rate across batch sizes merely fluctuates within a small range, unable to effectively explain the dramatic changes in speedup.
We further compare the behaviors of MoE and dense models in SD. We use Opt-30B, speculated by Opt-350M zhang2022opt , as the dense counterpart. To exclude the interference of factors besides model architecture and workload, we compare their target efficiency. As shown by Fig. 2(c), the target efficiency for the dense model continuously decreases. Although SD for MoE underperforms in small batches, it demonstrates greater potential across a wider range of larger batch sizes.
@@ 4.2 Impact of MoE Sparsity and Validation of Modeling Method
Figure 3: Comparison between GPU results and Modeling for MoEs with varying sparsity $\rho$ and $\gamma$.
To evaluate MoE sparsity’s impact on SD acceleration, we varied the number of activated experts per token ($K$). Directly changing $K$ without training affects the target model’s performance and speculation accuracy, so we adjust the speedup by multiplying the raw speedup with $\frac{\sigma_{K=8}}{\sigma_{K}}$, whose rationale is exhibited by Eq. 4. Fig. 3 shows the adjusted speedup alongside our modeling results for comparison. The parameters used in the modeling are decided using 21 GPU measurements, as explained in Section 3.3. The impact of measurement selection for parameter fitting on the modeling’s reliability is supplemented in the Appendix B.
There are three key observations. First, our modeling consistently aligns with experiment results across varying sparsity ($\rho$) and draft length ($\gamma$), validating our modeling’s reliability. Second, while the SD speedup in most MoEs exhibits an initial increase followed by a decrease, very sparse MoEs ($K$=1,2) show continuously decreasing speedup. This appears to conflict with the theoretical analysis, but after examining the components of our modeling, we identified the reason as follows. These very sparse MoEs have a disproportionately low ratio of FFN, thus making the memory-boundness of MoE FFN hard to manifest systematically as indicated by Amdahl’s Law. The Qwen2-57B-A14B model is designed based on $K$=8, but by reducing $K$ to 1 or 2, we actually artificially synthesize a model where Attention dominates. In practice, however, sparser MoEs typically incorporate more FFN parameters to maintain a balanced ratio between FFN and Attention components, resulting in acceleration patterns more similar to $K$=8 cases. Finally, as MoE models become sparser, the system’s transition from memory-bound to compute-bound is delayed. This is evidenced by two phenomena in the figure: With smaller $\rho$, (1) the batch size corresponding to peak speedup becomes larger; (2) the range of batch sizes that maintain speedup above a certain decay threshold (annotated by the brown dashed line in Fig. 3 for $\mathbf{x}/\sqrt{2}$) is wider. This validates our theoretical analysis and indicates that SD has broader applicability in sparser MoEs.
@@ 5 Conclusion and Limitation
Appendix A Proof of $\overline{T_{exp}}(T;\rho)$’s Trend with Varying $\rho$
In Section 3.2, Figure 1(c) demonstrates that: Given input token count $t=T>1$, the number of tokens each expert processes on average $\overline{T_{\text{exp}}}(T;\rho)=\frac{\rho T}{1-(1-\rho)^{T}}$ decreases as $\rho$ decreases. We prove this by showing $\frac{\mathrm{d}(\overline{T_{\text{exp}}}(T;\rho))}{\mathrm{d}\rho}>0$ when $T>1$.
@@ Appendix A Proof of $\overline{T_{exp}}(T;\rho)$’s Trend with Varying $\rho$
Appendix B The Impact of Measurement Selection for Parameter Fitting on Modeling
As mentioned in Section 3.3, since GPU execution is dynamic in practice, and not all operators are optimized to their theoretical limits, we introduced several parameters for relaxation. The values of these parameters are then automatically determined by fitting to GPU measurements under the least squares criterion. The results shown in Figure 3 in Section 4.2 were obtained by fitting to 21 GPU measurements. In this Section, we further explain:
  • Fitting Details of the modeling shown in Figure  3. (Section  B.1)
  • How the modeling is affected by alternative measurement selection. (Section  B.2)
@@ Appendix B The Impact of Measurement Selection for Parameter Fitting on Modeling
B.1 Fitting Details of the Modeling shown in Figure 3
@@
  • k1: It adjusts the intensity of the roofline effect of dense components. It should be larger than 0 to ensure the execution time increases as the token count increases. We don’t set a definite upper limit for $k1$, as its value is affected by other parameters. Given the hardware with fixed arithmetic units, the execution time grows linearly with the token count in the compute-bound regime. As shown in line 6 of Algorithm 1, $k1$ appears as a coefficient in the term $k1\cdot G(t;\lambda,s)$, whose gradient (representing the growing speed) in the compute-bound regime is $k1\cdot ln(s)\cdot s^{\lambda RP}$. As $s$ approaches 1, $k1$ needs to continuously increase to counterbalance $ln(s)$ that approaches 0. In summary, $k1_{min}=0$ and $k1_{max}=inf$.

### v3

@@ 4 Experiments
We conducted experiments on two pairs of MoE target models and draft models: Qwen2-57B-A14B-Instruct with Qwen2-0.5B-Instruct qwen2 , and Mixtral-8x7B-Instruct-v0.1 mixtral with Eagle speculation head li2024eagle . They represent two prevalent SD approaches: Qwen2 employs a standalone small model from the same model family as the draft model, while Mixtral uses a specifically trained speculation head. When we need to examine MoEs with different sparsity, we modify the num_experts_per_token in the model’s config.json file. For comparison with dense models, we use Opt-30b and Opt-350m zhang2022opt as the target and draft models. Models are evaluated on HumanEval humaneval and MT-bench mtbench datasets for code generation and conversation tasks, following previous works li2024eagle ; hass ; cai2401medusa . The tokenized prompt lengths range from 38 to 391 tokens for HumanEval and 5 to 356 tokens for MT-bench.
@@
We used the existing vllm kwon2023vllm framework for our experiments to verify theoretical predictions. Vllm supports batched speculative decoding, cudagraph optimization, and reports comprehensive data such as $_{D}, TT$$_{T}$, $T_{reject}$ and $\sigma$, thus being suitable for our experiments. To prevent unstable performance at the beginning, all data were obtained by averaging the results from the last five of the total ten runs. We conducted experiments on different hardware platforms including 2xGPU-A, 2xGPU-B, 4xGPU-A, 4xGPU-C.
@@ 4.1 Speedup Trend of Speculative Decoding for MoE
Figure 2: SD speedup (left y-axis) as a function of batch size and corresponding target efficiency values (right y-axis). Across different hardware platforms and MoE models, SD speedup first increases and then decreases, verifying our theoretical predictions. The target efficiency shows consistent trends with final speedup, validating its effectiveness.
[…]
@@
Mixtral humaneval 0.0 20.86 12.47 0.78 1.67 21.00 12.46 0.66 1.69 20.86 11.69 0.58 1.79
humaneval 1.0 21.52 15.58 0.61 1.38 21.39 16.03 0.46 1.33 21.48 16.23 0.39 1.32
mtbench 0.0 21.61 16.10 0.61 1.34 21.61 16.43 0.46 1.32 21.36 16.89 0.39 1.26
mtbench 1.0 21.33 17.70 0.53 1.21 21.33 17.84 0.43 1.20 21.33 18.05 0.35 1.18
[…]
Device Dataset Temp $\gamma=2$ $\gamma=3$ $\gamma=4$
2xGPU-B humaneval 0.0 15.96 9.34 0.95 1.71 15.96 7.95 0.93 2.01 15.96 6.96 0.90 2.29
@@
4xGPU-A humaneval 0.0 11.20 6.77 0.95 1.65 11.20 5.89 0.93 1.90 11.20 5.38 0.90 2.08
humaneval 1.0 11.72 8.51 0.81 1.38 12.05 8.30 0.73 1.45 11.23 7.70 0.67 1.46
mtbench 0.0 11.26 8.92 0.72 1.26 11.26 9.10 0.61 1.24 11.26 9.82 0.52 1.15
mtbench 1.0 11.78 10.32 0.67 1.14 11.30 9.42 0.58 1.20 11.78 11.25 0.47 1.05
4xGPU-C humaneval 0.0 17.84 10.00 0.95 1.79 17.84 8.33 0.93 2.14 17.84 7.94 0.90 2.25
humaneval 1.0 17.89 12.27 0.80 1.46 17.89 11.07 0.74 1.62 17.89 10.91 0.65 1.64
mtbench 0.0 20.40 15.87 0.71 1.29 20.40 16.22 0.62 1.26 20.40 16.33 0.54 1.25
mtbench 1.0 20.58 16.02 0.68 1.28 18.11 14.75 0.54 1.23 18.11 15.54 0.48 1.17
Figure 2 plots the end-to-end SD speedup (left y-axis) for MoE across various settings, validating our theoretical prediction about acceleration behavior. As batch size grows, speedup initially increases due to expert loading saturation, and then decreases due to compute-boundness. We denote the maximal speedup across batch sizes as $\mathbf{x}$ and summarize the results in Table 1. For both models, SD achieves higher acceleration with longer $\gamma$ for tasks with more predictable patterns (e.g., code generation) or less randomness (e.g., lower temperature), aligning with conclusions from previous research. Figure 5 in Appendix A.1 further presents SD speedup trends under more settings, including individual runs and their mean to show the statistical significance of our findings.
We further evaluate Qwen2-57B-A14B-Instruct on multiple hardware platforms (Table 2) to verify the generality of our conclusions. Combined with results of Qwen2 in Table 1, two observations can be made: (1) GPUs with higher ridge points yield larger SD speedups (e.g., 2×GPU-A vs. 2×GPU-B, 4×GPU-A vs. 4×GPU-C), since they provide more arithmetic units for verification. (2) Scaling from 2×GPU-A to 4×GPU-A reduces absolute runtimes ($T_{AR}$ and $T_{SD}$), but the SD speedups slightly degrade. This is because the large model benefits from inter-GPU parallelization, whereas the small draft model remains single-GPU, making its relative forward cost higher.
Figure 3: Comparison of target efficiency: MoE vs dense model.
Figure 2 also highlights the effectiveness of our metric target efficiency. It is computed as $\frac{T_{T}(B,1)}{T_{T}(B,\gamma)}$ as explained in Sec. 3.1, where both $T_{T}(B,1)$ and $T_{T}(B,\gamma)$ are obtained from vllm runtime logs. Target efficiency values are annotated on the right y-axis, showing a consistent trend with the end-to-end speedup. In contrast, the acceptance rate across batch sizes merely fluctuates within a small range, unable to effectively explain the dramatic changes in speedup.
We further compare the behaviors of MoE and dense models in SD. Since the effectiveness of target efficiency has been established in analyses above, and to avoid interference from algorithmic factors such as acceptance rate, we compare their target efficiency. As shown in Figure 3, the target efficiency for MoE first increases and then decreases, while that for the dense model decreases continuously. Consequently, although SD for MoE is less effective with small batches, it exhibits stronger potential across a wider range of larger batch sizes. Regarding end-to-end performance, SD speedups become more pronounced for MoE when the batch size exceeds 16, as supplemented in Figure 6 in Appendix A.2.
@@ 4.2 Impact of MoE Sparsity and Validation of Modeling Method
To evaluate MoE sparsity’s impact on SD acceleration, we varied the number of activated experts per token ($K$) of Qwen2-57B-A14B-Instruct. Directly changing $K$ without training affects the target model’s performance and speculation accuracy, so we adjust the speedup by multiplying the raw speedup with $\frac{\sigma_{K=8}}{\sigma_{K}}$, whose rationale is exhibited by Eq. 4. Fig. 4 shows the adjusted speedup alongside our modeling results for comparison. The parameters used in the modeling are decided using 21 GPU measurements, as explained in Section 3.3. The impact of measurement selection for parameter fitting on the modeling’s reliability is supplemented in Appendix C.
There are three key observations. First, the modeling consistently aligns with experiment results across varying sparsity ($\rho$) and draft length ($\gamma$), validating our modeling’s reliability.
Second, while the SD speedup in most MoEs exhibits an initial increase followed by a decrease, very sparse MoEs ($K=1,2$) show continuously decreasing speedup. This appears to conflict with the theoretical analysis, but after examining the components of our modeling, we identified the reason as follows. These very sparse MoEs have a disproportionately low ratio of FFN, thus making the memory-boundness of MoE FFN hard to manifest systematically as indicated by Amdahl’s Law. The Qwen2-57B-A14B model is designed based on $K=8$, but by reducing $K$ to 1 or 2, we actually artificially synthesize a model where Attention dominates. In practice, however, sparser MoEs typically incorporate more FFN parameters to maintain a balanced ratio between FFN and Attention components, resulting in acceleration patterns more similar to $K=8$ cases.
Finally, as MoE models become sparser, the system’s transition from memory-bound to compute-bound is delayed. This is evidenced by two phenomena in Fig. 4: With smaller $\rho$, (1) the batch size for the maximal speedup ($\mathbf{x}$) becomes larger; (2) the range of batch sizes that maintain speedup above a certain decay threshold (annotated by the brown dashed line in Fig. 4 for $\mathbf{x}/\sqrt{2}$) is wider. These validate our theoretical analysis and indicate that SD has broader applicability in sparser MoEs.
Figure 4: Comparison between GPU results and our modeling for Qwen2-57B-A14B-Instruct with varying sparsity $\rho$ and draft length $\gamma$.
@@ 5 Conclusion and Limitation
Acknowledgments and Disclosure of Funding
This work is supported by the National Natural Science Foundation of China (Grant Nos. 92267203) and Deng Feng Fund.
Appendix A Supplementary Experimental Results
This section presents additional experimental results referenced in Section 4, which are included here due to space limitations.
A.1 Trends of SD speedup under more configurations
Figure 5: SD speedup trends across more settings with individual runs and averages shown.
Figure 5 presents additional trends of SD speedup across different datasets, temperatures, and model types, serving as a supplement to Figure 2. The results demonstrate that SD speedup exhibits a consistent first-increase-then-decrease pattern, which aligns well with our theoretical analysis.
To confirm the statistical significance of our findings, we also present the five individual runs that constitute the averages in Figure 5. The variance across different runs is minimal, which is expected since the random seed is fixed across all runs to ensure identical workloads.
While the overall trend follows the first-increase-then-decrease pattern, local fluctuations are observable in the curves. For instance, Figure 5(c) exhibits a sawtooth-like decreasing trend. This phenomenon can be attributed to GPU quantization effects, as documented in NVIDIA’s documentation [47]. When dimensions are not evenly divisible by the GPU’s native tile sizes, computational performance degrades. AR decoding is more sensitive to this effect than SD, making the time ratio of AR to SD (namely, SD speedup) fluctuate. Despite these local variations, the overall speedup trend follows our theoretical predictions, confirming the validity of our conclusions.
A.2 End-to-end speedup comparison between MoE and dense models
Figure 6: End-to-end speedup comparison of MoE and dense models under various settings.
To isolate the effects of acceptance rate variations and enable a clearer focus on system bottlenecks, we have compared MoE and dense models using target efficiency in Section 4.1. In this section, we further compare their end-to-end speedup across various settings in Figure 6 as a supplement.
Two key observations emerge from the experiment results. First, while SD speedups for MoE models initially increase before declining, SD speedups for dense models continue to decrease. Consequently, SD achieves more substantial end-to-end speedups for MoE models at moderate batch sizes, which aligns with the trend in Figure 3 in Section 4.1. Second, the extent to which SD favors MoE over dense models varies across different configurations. For instance, at temperature = 1 (second row), SD demonstrates greater relative benefits for MoE compared to temperature = 0 (first row). This variation stems from diverse acceptance rates under different settings, which can obscure the observation of systemic bottlenecks. In summary, target efficiency serves as a reliable comparison metric while controlling for the confounding effects of algorithmic optimizations.
Appendix B Proof of $\overline{T_{exp}}(T;\rho)$’s Trend with Varying $\rho$
In Section 3.2, Fig. 1(c) demonstrates that: Given input token count $t=T>1$, the number of tokens each expert processes on average $\overline{T_{\text{exp}}}(T;\rho)=\frac{\rho T}{1-(1-\rho)^{T}}$ decreases as $\rho$ decreases. We prove this by showing $\frac{\mathrm{d}(\overline{T_{\text{exp}}}(T;\rho))}{\mathrm{d}\rho}>0$ when $T>1$.
@@ Appendix B Proof of $\overline{T_{exp}}(T;\rho)$’s Trend with Varying $\rho$
Appendix C More details about the Modeling Method
The main design considerations and expressions of our modeling method have been presented in Section 3.3. In this section, we provide additional content on the following topics to give a more comprehensive view of the modeling method:
  • Description and an illustrative diagram of the modeling process. (Appendix C.1)
  • Fitting Details of the modeling shown in Figure 4 in Section 4.1. (Appendix  C.2)
  • How the modeling is affected by alternative measurement selection. (Appendix  C.3)
@@ Appendix C More details about the Modeling Method
C.1 Description and Overview of the Modeling Process
Figure 7: The overall diagram of the modeling method.
Figure 7 presents the overall diagram of our modeling method. Building upon the theoretical analyses in Sections 3.1 and 3.2, we derive an expression for SD speedup as a function of workloads. This expression contains several relaxation parameters to be determined for complete modeling. We determine these parameters through empirical profiling. We first collect a small set of real measurements comprising various workloads and their corresponding SD speedups. We then perform parameter fitting using these measurements under the least squares criterion to obtain optimal parameter values. Since our theoretically-derived SD speedup expression already captures the fundamental performance tradeoffs, the fitting process is computationally lightweight and robust, as will be demonstrated in Appendix C.2 and C.3. Once the optimal parameters are obtained, the resulting expression can predict SD speedups for arbitrary workloads.
We now explain why profiling is necessary and why we cannot derive the complete SD speedup expression purely through theoretical analysis, examining both software and hardware considerations.
Software considerations: Actual execution times can deviate significantly from theoretical predictions for complex operators with diverse implementations. On GPUs, GEMM operations are indeed predictable due to their regular structure and highly optimized implementations. However, prediction becomes challenging for operators such as Attention, which involve customized kernel optimizations (e.g., FlashAttention1/2, eager attention, SDPA attention) and operator fusion strategies (incorporating various nonlinear layers or positional encodings such as RoPE and its variants). To illustrate this complexity, we examine profiling results from Qwen2-57B-A14B (hidden size 3584) and Mixtral-8x7B (hidden size 4096). For FFN, Qwen takes a shorter time than Mixtral (143us vs 226us), aligning with their relative hidden sizes. However, for Attention, Qwen takes a longer time than Mixtral (271us vs 115us), contradicting the theoretical expectation based on hidden size scaling.
Hardware considerations: GPU microarchitectures vary across different series, which can greatly impact execution times. For instance, attention efficiency depends heavily on hardware-aware programming optimizations, while different GPUs vary in cache configurations and thread-memory interaction patterns across memory hierarchy levels. By taking advantage of new capabilities in modern hardware, FlashAttention-3 successfully increases GPU utilization from 35% to 75% on H100 GPUs [48]. Moreover, many critical hardware details remain undisclosed by GPU vendors, making theoretical predictions impractical.
Therefore, modeling speedup trends with pure analytical methods requires case-by-case analysis for different operator implementations and GPU microarchitectures. In contrast, the hyperparameter approach offers a more generalizable paradigm and is easy to use: all parameters possess clear physical interpretations, only minimal profiling data are required, and the computational overhead is low. Our method achieves a balance between effectiveness and practicality: on one hand, it captures the primary performance drivers (i.e. the number of activated experts and roofline trends); on the other hand, it avoids getting entangled in low-level implementation complexities. This approach is also used by other system optimization frameworks such as NanoFlow [49], which similarly adopt a two-stage strategy of profiling followed by runtime execution.
C.2 Fitting Details of the Modeling shown in Figure 4
@@
  • k1: It adjusts the intensity of the roofline effect of dense components. It should be larger than 0 to ensure the execution time increases as the token count increases. We don’t set a definite upper limit for $k1$, as its value is affected by other parameters. Given the hardware with fixed arithmetic units, the execution time grows linearly with the token count in the compute-bound regime. As shown in line 6 of Algorithm 1, $k1$ appears as a coefficient in the term $k1\cdot G(t;\lambda,s)$, whose gradient in the compute-bound regime is $k1\cdot ln(s)\cdot s^{\lambda RP}$. As $s$ approaches 1, $k1$ needs to continuously increase to counterbalance $ln(s)$ that approaches 0.
  • k2: It represents the time required to load one expert. Given a target model, we denote the parameter count per expert as $V_{exp}$. Consequently, the theoretical minimum loading time can be calculated as $\text{{k2}}_{min}=\frac{V_{exp}\times bitwidth}{\text{{peak memory bandwidth}}}$. For the upper bound of the relaxation range, we set $\text{{k2}}_{max}=5\times k_{2min}$.
  • k3: It adjusts the intensity of the roofline effect of sparse components. Similar to k1, we set $\text{{k3}}_{min}=0$ and $\text{{k3}}_{max}=\text{{inf}}$.
  • draft_bias: It represents the time required to load the dense draft model. We denote the draft model’s parameter count as $V_{draft}$. Consequently, the theoretical minimum loading time can be calculated as $\text{{draft\_bias}}_{min}=\frac{V_{draft}\times bitwidth}{\text{{peak memory bandwidth}}}$. For the upper bound of the relaxation range, we set $\text{{draft\_bias}}_{max}=5\times\text{{draft\_bias}}_{min}$.
  • draft_k: It adjusts the intensity of the roofline effect of the dense draft model. Similar to k1, we set $\text{{draft\_k}}_{min}=0$ and $\text{{draft\_k}}_{max}=\text{{inf}}$.
  • reject_bias: It represents the fixed overhead when performing rejection sampling. Vllm reports its elapsed time during SD, and we denote the maximum across measurements as $T_{rej}$. We then set $\text{{reject\_bias}}_{min}=0$ and $\text{{reject\_bias}}_{max}=T_{rej}$.
  • reject_k: It represents the incremental processing time in rejection sampling as the input token count increases. For simplicity, we set $\text{{reject\_k}}_{min}=0$ and $\text{{reject\_k}}_{max}=T_{rej}$ just like reject_bias.
  • $\lambda$: It represents the ratio of the empirical ridge point to the theoretical ridge point. Since memory bandwidth is typically less utilized than arithmetic units, we set $\lambda_{min}=0.2$ and $\lambda_{max}=1$.
C.3 Exploration of Alternative Measurement Selection
In this section, we demonstrate the impact of varying the number ($m$) of measurements used for fitting on the modeling results. Given that our model incorporates 10 parameters, a minimum of 10 profiling data points ($m\geq 10$) are required to determine all parameters. We present the modeling fitting with $m$ ranging from 10 to 228. The data selection method follows the stride-based approach described in the previous section, specifically M = df[begin:end:stridestride]. Measurement count $m$ and stridestride satisfy the following relation: $m=\lceil 228/stride\rceil$.